The mole fraction of solvent in 0.1 molal aqua solution is

Grade:12th pass

2 Answers

Ashish Kumar
6 years ago
Your question is a bit strange one but I’ll answer both the possibilities.One is yours and the other too which i assume is the correct answer.
1.00 molal aqueous solution means that 1 mol of solute is dissolved in 1 Kg. of water.
Number of moles of solute = 1
Number of moles of water = 1000/18= 55.556
mole fraction of solvent =
Moles of solvent/moles of solute + moles of solution
=>55.56/1+55.56=0.98 Answer
Now  the question which I assume is the correct one is-“The mole fraction of solute in 0.1 molal aqua solution is”
so for this, mole fraction= moles of solute/moles of solute+moles of solvent =0.1768 Answer
 
Ashish Kumar
6 years ago
Ohh..man i didn’t realise that it was 0.1....sorry for the mistake so posting the correct answer-
0.1 molal aqueous solution means that 1 mol of solute is dissolved in 10 Kg. of water.
Number of moles of solute = 1
Number of moles of water = 10000/18= 555.56
mole fraction of solvent =
Moles of solvent/moles of solute + moles of solution
=>555.6/1+555.6=0.998 Answer
Now  the question which I assume is the correct one is-“The mole fraction of solute in 0.1 molal aqua solution is”
so for this, mole fraction= moles of solute/moles of solute+moles of solvent =0.001796 Answer
 
Apologies.....please now if I solved your query do approve my answer...thank you!